# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 9.4 Absorption and edges
The first major goal of this subsection is to prove the connection between absorption and stability under edges stated in Theorems 5.19 and 5.21. The following observation will be often implicitly used in proofs: if $B$ is stable under semilattice (majority, abelian) edges in an algebra $\mathbf{A}$ and $\mathbf{C}\leq\mathbf{A}$, then $B\cap C$ is stable under semilattice (majority, abeialn) edges in the algebra $\mathbf{C}$.
@@
Let $\mathbf{A}$ be a minimal Taylor algebra and $B\subseteq A$. If $B$ absorbs $\mathbf{A}$, then $B$ is stable under all the abelian and semilattice edges. Moreover, if $B$ 2-absorbs $\mathbf{A}$, then $B$ is stable under all the edges.
Proof.
Assume that $B$ absorbs $\mathbf{A}$ and $(b,a)$ is an edge with witnessing congruence $\theta$ of $\mathbf{E}:=\operatorname{Sg}_{\mathbf{A}}(b,a)$ such that $b/\theta$ intersects $B$. By Theorem 5.5, $B$ is a subuniverse of $\mathbf{A}$. As $B\trianglelefteq\mathbf{A}$ we also have $B\cap E\trianglelefteq\mathbf{E}$ and then $b/\theta\in(B\cap E)/\theta\trianglelefteq\mathbf{E}/\theta$.
By Theorem 5.13, if $(b,a)$ is a semilattice edge, then $\mathbf{E}/\theta$ is term equivalent to the two element semilattice with absorbing element $a/\theta$. By the description of term operations in the two element semilattice (see Subsection 2.2), the only nontrivial absorbing subuniverse of $\mathbf{E}/\theta$ is $\{a/\theta\}$, therefore $(B\cap E)/\theta=E/\theta$, so each $\theta$-class indeed intersects $B$. If $(b,a)$ is an abelian edge, then $\mathbf{E}/\theta$ is term equivalent to an affine Mal’cev algebra. These algebras do not have any nontrivial absorbing subuniverses (since, e.g., the term operations are of the form $\sum a_{i}x_{i}$ where $\sum a_{i}=1$ and these operations do not witness any nontrivial absorption) and we get $(B\cap E)/\theta=E/\theta$ in this case as well. If, additionally, $B$ 2-absorbs $\mathbf{A}$, then $(B\cap E)/\theta\trianglelefteq_{2}\mathbf{E}/\theta$ and we get the same conclusion by the description of term operations in the two-element majority algebra. ∎
@@
Let $\mathbf{A}$ be a minimal Taylor algebra.
@@
Recall (Proposition 5.15) that for an s-edge $(a,b)$ in a minimal Taylor algebra $\mathbf{A}$, the set $\{a,b\}$ is a subuniverse of $\mathbf{A}$ (so the witnessing congruence for this edge is the equality relation on $\{a,b\}$). It follows that every set stable under semilattice edges is s-closed.
@@
Suppose that $B$ is an s-closed subset of a minimal Taylor algebra $\mathbf{A}$. Then the following are equivalent:
@@
    for all $a\in A\setminus B$ and all $b\in B$, the algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)$ has a nontrivial 2-absorbing subuniverse,
@@
    for all $a\in A$ and all $b\in B$, there is a directed s-walk from $a$ to an element in $B$ which is contained in $\operatorname{Sg}_{\mathbf{A}}(a,b)$.
Proof.
To see that (a) implies (b), note that $B\trianglelefteq_{2}\mathbf{A}$ implies that $B\cap\operatorname{Sg}_{\mathbf{A}}(a,b)\trianglelefteq_{2}\operatorname{Sg}_{\mathbf{A}}(a,b)$.
We prove that (b) implies (c) by induction on the size of $\mathbf{A}$ (assuming that the implication is true for all algebras of strictly smaller size). Let $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$ and let $C$ be a nontrivial 2-absorbing subuniverse of $\mathbf{E}$. By Proposition 9.7 there is an s-walk from $a$ to some $a^{\prime}\in C$ contained in $E$: either $a\not\in C$ and there is an s-edge $(a,a^{\prime})$ by that proposition, or $a\in C$ and we have the trivial s-walk $a$ from $a$ to $a$. If $a^{\prime}\in B$, then we are done, so we assume $a^{\prime}\not\in B$. Similarly, there is an s-walk from $b$ to some $b^{\prime}\in C$ and, since $B$ is s-closed, we have $b^{\prime}\in B$. Let $\mathbf{A}^{\prime}=\operatorname{Sg}(a^{\prime},b^{\prime})$ and $B^{\prime}=B\cap A^{\prime}$. Note that $\mathbf{A}^{\prime}$ is a proper subalgebra of $\mathbf{A}$ (since $A^{\prime}\subseteq C\subsetneq E\subseteq A$), that $B^{\prime}$ is an s-closed subset of $\mathbf{A}^{\prime}$ (since $B$ is s-closed in $\mathbf{A}$ and $B^{\prime}\subseteq B$), and that $\mathbf{A}^{\prime}$ with $B^{\prime}$ satisfies the condition of item (b). By induction hypothesis there is an s-walk from $a^{\prime}$ to some element $b^{\prime\prime}\in B^{\prime}$ contained in $\operatorname{Sg}(a^{\prime},b^{\prime\prime})\subseteq E$. But we also have a walk from $a$ to $a^{\prime}$ contained in $E$ and we are done by concatenating these two walks.
Now suppose that (c) holds. Let $\mathbf{F}$ be the subalgebra of $\mathbf{A}^{A^{2}}$ formed by all the binary term operations of $\mathbf{A}$ (recall Subsection 7.4) and let $\mathbf{R}\leq\mathbf{A}^{X}$ be the projection of $\mathbf{F}$ onto the set of coordinates $X=\{(a,b),(b,a):a\in A,b\in B\}$. Notice that any $f\in R\cap B^{X}$ witnesses that $B$ is a 2-absorbing subset of $\mathbf{A}$, therefore it is enough to show that $R$ intersects $B^{X}$. Assume that this is not the case, for the sake of proving a contradiction.
@@
Note that the definition in the second item makes sense since $\operatorname{proj}_{{\mathbf{x}}}R$ contains an element of $B$ (as witnessed by one of the two projection operations in $R$), so such an s-walk $a_{1},\dots,a_{n}$ exists by (c). Now $\{a_{2}\}$ is a 2-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a_{1},a_{2})=\{a_{1},a_{2}\}$ since $(a_{1},a_{2})$ is an s-edge. Therefore the subuniverse $R^{\prime}$ of $\mathbf{R}$ obtained by fixing the coordinate ${\mathbf{x}}$ of $R$ to $\{a_{2}\}$ 2-absorbs the subalgebra of $\mathbf{R}$ obtained by fixing the same coordinate to $\{a_{1},a_{2}\}$. By Proposition 9.7, there exists an s-edge from $f$ to an element of $R^{\prime}$, giving us an s-edge $(f,g)$ in $R$ for some $g\in R$ with $g({\mathbf{x}})=a_{2}$. Since $(f,g)$ is an s-edge, the pair $(f({\mathbf{y}}),g({\mathbf{y}}))$ is, for any ${\mathbf{y}}\in Y$, also an s-edge (witnessed by the same term). Recalling that $B$ is s-closed we obtain $g|_{Y}\in B^{Y}$. But $g({\mathbf{x}})$ starts the s-walk $a_{2},\dots,a_{n}$, which is one step shorter than the original walk $a_{1},\dots,a_{n}$, a contradiction to the minimality of $n$ (if $n>2$) or the maximality of $Y$ (if $n=2$). ∎
@@
Proof.
Factoring by a maximal congruence we can assume that $\mathbf{A}$ is simple (as 3-absorbing subuniverses lift from quotients). If $\mathbf{A}$ has a ternary strongly functional subpower or a binary irredundant subdirect subpower, then we are done by Proposition 4.8 or Corollary 4.5 (since the subpower is then necessarily linked by the simplicity of $\mathbf{A}$). Also, if $(a,b)$ is an s-edge, we are done as well.
Otherwise, $\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$ is not full (it does not contain $(b,b)$), so it must be a graph of a bijection – an automorphism of $\mathbf{A}$. Moreover, by Theorem 4.7, $\mathbf{A}$ has no proper irredundant subdirect subpowers, therefore $\operatorname{Sg}_{\mathbf{A}^{3}}((b,a,a),(a,b,a),(a,a,b))$ is full. In particular, it contains $(a,a,a)$, therefore there exists a term operation $t$ such $t(b,a,a)=t(a,b,a)=t(a,a,b)=a$. Since $\mathbf{A}$ has an automorphism switching $a$ and $b$ we see that $t$ acts as a majority operation on $\{a,b\}$. Now $\{a,b\}$ is a subuniverse of $\mathbf{A}$ by Proposition 5.3, therefore $A=\{a,b\}$ as $\mathbf{A}$ is generated by $a$,$b$. The majority operation witnesses that $\{a\}$ (and $\{b\}$) 3-absorbs $\mathbf{A}$. ∎
@@
If $\mathbf{A}$ has a nontrivial affine quotient, then since $B$ is stable under abelian edges, $B$ must intersect the congruence class which contains $a$, so we can apply the inductive assumption to find an s-edge from $a$ to $B$ which is contained in that congruence class. Otherwise, Lemma 9.11 implies that there is some nontrivial 3-absorbing subalgebra $\mathbf{C}$ of $\mathbf{A}$.
Next we show that $C$ intersects $B$. Suppose the contrary and apply Proposition 9.7 to the element $b$ and $C\trianglelefteq_{3}\mathbf{A}$ – we get an s-edge from $b$ to $C$, which is impossible since $B$ is s-closed, or a majority edge $(b,c)$ with $c\in C$ such that the witnessing congruence $\theta$ of $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(b,c)$ has two equivalence classes, $C\cap E$ and $E\setminus C$, which is impossible since $B$ is stable under majority edges. (Note that we can get a stronger form of instability in the majority case as follows. We get $B\cap E\trianglelefteq_{2}E\setminus C$ by induction hypothesis, therefore $B\cap E\trianglelefteq_{3}\mathbf{E}$ by transitivity of 3-absorption stated in item (3) of Proposition 5.12. Then the union of $B\cap E$ and $C\cap E$ is a subuniverse of $\mathbf{E}$ by item (1) in Proposition 5.12. Since $E$ is generated by $b\in B\cap E$ and $c\in C\cap E$, we obtain $E\setminus C=B\cap E$.)
Now we have $C\cap B\neq\emptyset$ and we can apply the induction hypothesis to see that $C\cap B$ 2-absorbs $\mathbf{C}$. This implies that $C\cap B$ is a 3-absorbing subalgebra of $\mathbf{A}$ by the transitivity of 3-absorption (item (3) in Proposition 5.12). Thus we may assume without loss of generality that $C\subseteq B$.
By Proposition 9.7, there is either a semilattice edge from $a$ to $c\in C$, or a majority edge from $a$ to $c\in C$ witnessed by the congruence on $\operatorname{Sg}(a,c)$ with classes $\operatorname{Sg}(a,c)\cap C$ and $\operatorname{Sg}(a,c)\setminus C$. In the first case we are done, so suppose we are in the second case. Since $\operatorname{Sg}(a,c)\cap C\subseteq B$ and $B$ is stable under majority edges, the congruence class $\operatorname{Sg}(a,c)\setminus C$ has a nonempty intersection with $B$. Then we may apply the induction hypothesis to $\operatorname{Sg}(a,c)\setminus C$ to see that there is a semilattice edge from $a$ to $(\operatorname{Sg}(a,c)\setminus C)\cap B$, which finishes the proof. ∎
@@
Let $\mathbf{A}$ be a simple algebra and $R\leq_{sd}\mathbf{A}^{3}$ be a symmetric ternary relation whose projection to each pair of coordinates is full. For any $a\in A$ denote by $R_{a}$ the relation $R_{a}(x,y)\equiv R(a,x,y)$. Then either $\mathbf{A}$ is abelian or the set $B=\{a\in A:R_{a}\mbox{ is linked }\}$ is a nonempty strongly projective subuniverse of $\mathbf{A}$.
Proof.
Consider the ternary relation $S$ defined by $S(x,y,z)$ if $y$ and $z$ are ”left-linked” in $R_{x}$, that is, $(y,z)$ is in the relation $(R_{x}-R_{x})+\cdots+(R_{x}-R_{x})$ with a sufficient number of summands. Note that $S$ is pp-definable from $R$, so $S$ is a subuniverse of $\mathbf{A}^{3}$. Since the projection of $R$ onto each pair of coordinates is full, the relation $R_{a}$ is subdirect for every $a$. If $R_{a}$ is linked, then $S_{a}$ (defined analogously to $R_{a}$) is full. If $R_{a}$ is not linked, then it is the graph of a bijection $A\to A$ (since $\mathbf{A}$ is simple), so $S_{a}$ is the equality relation. In summary, $S$ is equal to the relation $B(x)\vee(y=z)$. By Proposition 7.3, $B$ is strongly projective.
If $B$ is empty, then every $R_{a}$ is the graph of a bijection. But then, by symmetry of $R$, fixing any coordinate to any $a$ and projecting out this coordinate gives the graph of a bijection. Therefore $R$ is strongly functional, so $\mathbf{A}$ is abelian by Proposition 4.8, finishing the proof. ∎
We are ready to prove the main tool for Theorem 5.21. It will be useful to generalize the concept of a $B$-essential relation introduced in Subsection 7.5 to infinite powers: A subuniverse $R$ of $\mathbf{A}^{X}$ is $B$-essential if $R$ does not intersect $B^{X}$ but every projection of $R$ onto all but one of the coordinates intersects the corresponding power of $B$. By Proposition 7.2, $\{a\}$ absorbs $\mathbf{A}$ by a term operation of arity $n$ if and only if it does not have any $\{a\}$-essential subpower of arity $n$.
@@
Let $\mathbf{A}$ be a Taylor algebra and $a\in A$. Suppose that $\{a\}$ does not absorb $\mathbf{A}$ but does absorb every proper subalgebra of $\mathbf{A}$ that contains $\{a\}$. Then
  • $\mathbf{A}$ has a nontrivial abelian quotient or
  • there exists a nonempty projective subuniverse $B$ of $\mathbf{A}$ such that $a\not\in B$.
Proof.
@@
First, observe that it is enough to prove the claim for simple $\mathbf{A}$. Indeed, if $\alpha$ is a maximal congruence of $\mathbf{A}$, then either $a/\alpha$ absorbs $A$ (in which case $\{a\}$ absorbs $\mathbf{A}$ by transitivity of absorption, a contradiction to the assumptions), or it does not (in which case we apply the simple case – observe that projective subuniverses lift from quotients). Assume therefore that $\mathbf{A}$ is simple.
@@
We further assume that every relation $S\leq_{sd}\mathbf{A}^{2}$ whose left center contains $a$ is full. Indeed, if it is not, then by Corollary 8.5 either $\mathbf{A}$ has a proper projective and 2-absorbing subuniverse $B$ or the left center of $R$ is an absorbing subuniverse of $\mathbf{A}$. By transitivity of absorption and the assumptions, no absorbing subuniverse of $\mathbf{A}$ can contain $a$. Therefore $B$ would be a projective subuniverse not containing $a$ and the proof would be concluded.
@@
Next we observe that there exists a symmetric $\{a\}$-essential relation $R\leq\mathbf{A}^{\mathbb{N}}$ (here it will be convenient to use an infinitary relation). Indeed, for any arity $n$ there exists by Proposition 7.2 some $\{a\}$-essential relation, i.e., containing tuples $(b_{1},a,a,\dots,a)$, $(a,b_{2},a\dots,a),\dots$ and not containing $(a,a,\dots,a)$. Taking an element $b=b(n)$ that appears the most often among the $b_{i}$ and fixing (and projecting out) the other coordinates to $a$ we get an $\{a\}$-essential relation of arity at least $n/|A|$ containing $(b,a,a\dots)$, $(a,b,a,\dots)$, …. Finally, by taking $R=\operatorname{Sg}_{\mathbf{A}^{\mathbb{N}}}((b,a,a,\dots,),(a,b,a,\dots),\dots)$, where $b$ is an element that appears infinitely many times among the $b(n)$, we get a symmetric $\{a\}$-essential relation.
@@
For a symmetric relation $R\leq\mathbf{A}^{\mathbb{N}}$, denote $B(R)=\{b\in A:(b,a,a,\dots)\in R\}$, and take a symmetric $\{a\}$-essential relation $R$ (i.e., $a\not\in B(R)\neq\emptyset$) such that $B(R)$ is maximal. Set $B=B(R)$. We will show that $R$ either gives us a ternary subpower forcing strong projectivity or abelianess, or $R$ is a witness for $B$ being a projective subuniverse, i.e., after fixing any co-finite collection coordinates to $a$ and projecting them out, $R$ becomes $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{k})$ (see Proposition 2.6); observe here that the obtained relation is indeed a subpower of $\mathbf{A}$ despite the infinite arity of $R$. Since $a\not\in B$ we will be done.
@@
We claim that $\operatorname{Sg}_{\mathbf{A}}(a,b)=A$ for every $b\in B$. Indeed, otherwise $R$ defined by $\operatorname{Sg}_{\mathbf{A}^{\mathbb{N}}}\{(b,a,a,\dots)$, $(a,b,a,\dots),\dots\}$ is an $\{a\}$-essential symmetric subuniverse of $(\operatorname{Sg}_{\mathbf{A}}(a,b))^{\mathbb{N}}$, so $\{a\}$ does not absorb $\operatorname{Sg}_{\mathbf{A}}(a,b)$ by Proposition 7.2 (consider, as above, the relations obtained by fixing co-finite collections of coordinates to $a$), contradicting the assumptions of the theorem.
@@
If $B^{\prime}(=B+S+S)=B$, then $S$ is not linked, so it is a graph of a bijection by simplicity of $\mathbf{A}$. We fix all but three arbitrarily selected coordinates of $R$ to $a$, project onto the three coordinates, and call $T$ the obtained subuniverse of $\mathbf{A}^{3}$. The relation $T$ has full projection to the first two coordinates (as we argued above) and the binary relation obtained by fixing a coordinate to $a$ is the graph of a bijection. Lemma 9.12 now implies that either $\mathbf{A}$ is abelian or $\mathbf{A}$ has a nonempty strongly projective subuniverse that does not contain $a$. Since strong projectivity implies projectivity, we are done.

### v6

@@ 9.4 Absorption and edges
The first major goal of this subsection is to prove the connection between absorption and stability under edges stated in Theorems 5.19 and 5.21. The following observation will be often implicitly used in proofs: if $B$ is stable under semilattice (majority, abelian) edges in an algebra and , then $B\cap C$ is stable under semilattice (majority, abeialn) edges in the algebra .
@@
Let be a minimal Taylor algebra and $B\subseteq A$. If $B$ absorbs , then $B$ is stable under all the abelian and semilattice edges. Moreover, if $B$ 2-absorbs , then $B$ is stable under all the edges.
Proof 9.26.
Assume that $B$ absorbs and $(b,a)$ is an edge with witnessing congruence $\theta$ of such that $b/\theta$ intersects $B$. By Theorem 5.5, $B$ is a subuniverse of . As $B\trianglelefteq$ we also have $B\cap E\trianglelefteq$ and then $b/\theta\in(B\cap E)/\theta\trianglelefteq$.
By Theorem 5.13, if $(b,a)$ is a semilattice edge, then is term equivalent to the two element semilattice with absorbing element $a/\theta$. By the description of term operations in the two element semilattice (see Subsection 2.2), the only nontrivial absorbing subuniverse of is $\{a/\theta\}$, therefore $(B\cap E)/\theta=E/\theta$, so each $\theta$-class indeed intersects $B$. If $(b,a)$ is an abelian edge, then is term equivalent to an affine Mal’cev algebra. These algebras do not have any nontrivial absorbing subuniverses (since, e.g., the term operations are of the form $\sum a_{i}x_{i}$ where $\sum a_{i}=1$ and these operations do not witness any nontrivial absorption) and we get $(B\cap E)/\theta=E/\theta$ in this case as well. If, additionally, $B$ 2-absorbs , then $(B\cap E)/\theta\trianglelefteq_{2}$ and we get the same conclusion by the description of term operations in the two-element majority algebra.
@@
Let be a minimal Taylor algebra.
@@
Recall (Proposition 5.15) that for an s-edge $(a,b)$ in a minimal Taylor algebra , the set $\{a,b\}$ is a subuniverse of (so the witnessing congruence for this edge is the equality relation on $\{a,b\}$). It follows that every set stable under semilattice edges is s-closed.
@@
Suppose that $B$ is an s-closed subset of a minimal Taylor algebra . Then the following are equivalent:
@@
    for all $a\in A\setminus B$ and all $b\in B$, the algebra $\operatorname{Sg}(a,b)$ has a nontrivial 2-absorbing subuniverse,
@@
    for all $a\in A$ and all $b\in B$, there is a directed s-walk from $a$ to an element in $B$ which is contained in $\operatorname{Sg}(a,b)$.
Proof 9.29.
To see that (a) implies (b), note that $B\trianglelefteq_{2}$ implies that $B\cap\operatorname{Sg}(a,b)\trianglelefteq_{2}\operatorname{Sg}(a,b)$.
We prove that (b) implies (c) by induction on the size of (assuming that the implication is true for all algebras of strictly smaller size). Let and let $C$ be a nontrivial 2-absorbing subuniverse of . By Proposition 9.22 there is an s-walk from $a$ to some $a^{\prime}\in C$ contained in $E$: either $a\not\in C$ and there is an s-edge $(a,a^{\prime})$ by that proposition, or $a\in C$ and we have the trivial s-walk $a$ from $a$ to $a$. If $a^{\prime}\in B$, then we are done, so we assume $a^{\prime}\not\in B$. Similarly, there is an s-walk from $b$ to some $b^{\prime}\in C$ and, since $B$ is s-closed, we have $b^{\prime}\in B$. Let and $B^{\prime}=B\cap A^{\prime}$. Note that is a proper subalgebra of (since $A^{\prime}\subseteq C\subsetneq E\subseteq A$), that $B^{\prime}$ is an s-closed subset of (since $B$ is s-closed in and $B^{\prime}\subseteq B$), and that with $B^{\prime}$ satisfies the condition of item (b). By induction hypothesis there is an s-walk from $a^{\prime}$ to some element $b^{\prime\prime}\in B^{\prime}$ contained in $\operatorname{Sg}(a^{\prime},b^{\prime\prime})\subseteq E$. But we also have a walk from $a$ to $a^{\prime}$ contained in $E$ and we are done by concatenating these two walks.
Now suppose that (c) holds. Let be the subalgebra of formed by all the binary term operations of (recall Subsection 7.4) and let be the projection of onto the set of coordinates $X=\{(a,b),(b,a):a\in A,b\in B\}$. Notice that any $f\in R\cap B^{X}$ witnesses that $B$ is a 2-absorbing subset of , therefore it is enough to show that $R$ intersects $B^{X}$. Assume that this is not the case, for the sake of proving a contradiction.
@@
Note that the definition in the second item makes sense since $\operatorname{proj}_{{\mathbf{x}}}R$ contains an element of $B$ (as witnessed by one of the two projection operations in $R$), so such an s-walk $a_{1},\dots,a_{n}$ exists by (c). Now $\{a_{2}\}$ is a 2-absorbing subuniverse of $\operatorname{Sg}(a_{1},a_{2})=\{a_{1},a_{2}\}$ since $(a_{1},a_{2})$ is an s-edge. Therefore the subuniverse $R^{\prime}$ of obtained by fixing the coordinate ${\mathbf{x}}$ of $R$ to $\{a_{2}\}$ 2-absorbs the subalgebra of obtained by fixing the same coordinate to $\{a_{1},a_{2}\}$. By Proposition 9.22, there exists an s-edge from $f$ to an element of $R^{\prime}$, giving us an s-edge $(f,g)$ in $R$ for some $g\in R$ with $g({\mathbf{x}})=a_{2}$. Since $(f,g)$ is an s-edge, the pair $(f({\mathbf{y}}),g({\mathbf{y}}))$ is, for any ${\mathbf{y}}\in Y$, also an s-edge (witnessed by the same term). Recalling that $B$ is s-closed we obtain $g|_{Y}\in B^{Y}$. But $g({\mathbf{x}})$ starts the s-walk $a_{2},\dots,a_{n}$, which is one step shorter than the original walk $a_{1},\dots,a_{n}$, a contradiction to the minimality of $n$ (if $n>2$) or the maximality of $Y$ (if $n=2$).
@@
Proof 9.31.
Factoring by a maximal congruence we can assume that is simple (as 3-absorbing subuniverses lift from quotients). If has a ternary strongly functional subpower or a binary irredundant subdirect subpower, then we are done by Proposition 4.8 or Corollary 4.5 (since the subpower is then necessarily linked by the simplicity of ). Also, if $(a,b)$ is an s-edge, we are done as well.
Otherwise, $\operatorname{Sg}((a,b),(b,a))$ is not full (it does not contain $(b,b)$), so it must be a graph of a bijection – an automorphism of . Moreover, by Theorem 4.7, has no proper irredundant subdirect subpowers, therefore $\operatorname{Sg}((b,a,a),(a,b,a),(a,a,b))$ is full. In particular, it contains $(a,a,a)$, therefore there exists a term operation $t$ such $t(b,a,a)=t(a,b,a)=t(a,a,b)=a$. Since has an automorphism switching $a$ and $b$ we see that $t$ acts as a majority operation on $\{a,b\}$. Now $\{a,b\}$ is a subuniverse of by Proposition 5.3, therefore $A=\{a,b\}$ as is generated by $a$,$b$. The majority operation witnesses that $\{a\}$ (and $\{b\}$) 3-absorbs .
@@
If $\mathbf{A}$ has a nontrivial affine quotient, then since $B$ is stable under abelian edges, $B$ must intersect the congruence class which contains $a$, so we can apply the inductive assumption to find an s-edge from $a$ to $B$ which is contained in that congruence class. Otherwise, Lemma 9.30 implies that there is some nontrivial 3-absorbing subalgebra of $\mathbf{A}$.
Next we show that $C$ intersects $B$. Suppose the contrary and apply Proposition 9.22 to the element $b$ and $C\trianglelefteq_{3}$ – we get an s-edge from $b$ to $C$, which is impossible since $B$ is s-closed, or a majority edge $(b,c)$ with $c\in C$ such that the witnessing congruence $\theta$ of has two equivalence classes, $C\cap E$ and $E\setminus C$, which is impossible since $B$ is stable under majority edges. (Note that we can get a stronger form of instability in the majority case as follows. We get $B\cap E\trianglelefteq_{2}E\setminus C$ by induction hypothesis, therefore $B\cap E\trianglelefteq_{3}$ by transitivity of 3-absorption stated in item (3) of Proposition 5.12. Then the union of $B\cap E$ and $C\cap E$ is a subuniverse of  by item (1) in Proposition 5.12. Since $E$ is generated by $b\in B\cap E$ and $c\in C\cap E$, we obtain $E\setminus C=B\cap E$.)
Now we have $C\cap B\neq\emptyset$ and we can apply the induction hypothesis to see that $C\cap B$ 2-absorbs . This implies that $C\cap B$ is a 3-absorbing subalgebra of $\mathbf{A}$ by the transitivity of 3-absorption (item (3) in Proposition 5.12). Thus we may assume without loss of generality that $C\subseteq B$.
By Proposition 9.22, there is either a semilattice edge from $a$ to $c\in C$, or a majority edge from $a$ to $c\in C$ witnessed by the congruence on $\operatorname{Sg}(a,c)$ with classes $\operatorname{Sg}(a,c)\cap C$ and $\operatorname{Sg}(a,c)\setminus C$. In the first case we are done, so suppose we are in the second case. Since $\operatorname{Sg}(a,c)\cap C\subseteq B$ and $B$ is stable under majority edges, the congruence class $\operatorname{Sg}(a,c)\setminus C$ has a nonempty intersection with $B$. Then we may apply the induction hypothesis to $\operatorname{Sg}(a,c)\setminus C$ to see that there is a semilattice edge from $a$ to $(\operatorname{Sg}(a,c)\setminus C)\cap B$, which finishes the proof.
@@
Let be a simple algebra and $R\leq_{sd}$ be a symmetric ternary relation whose projection to each pair of coordinates is full. For any $a\in A$ denote by $R_{a}$ the relation $R_{a}(x,y)\equiv R(a,x,y)$. Then either is abelian or the set $B=\{a\in A:R_{a}\mbox{ is linked }\}$ is a nonempty strongly projective subuniverse of .
Proof 9.34.
Consider the ternary relation $S$ defined by $S(x,y,z)$ if $y$ and $z$ are ”left-linked” in $R_{x}$, that is, $(y,z)$ is in the relation $(R_{x}-R_{x})+\cdots+(R_{x}-R_{x})$ with a sufficient number of summands. Note that $S$ is pp-definable from $R$, so $S$ is a subuniverse of . Since the projection of $R$ onto each pair of coordinates is full, the relation $R_{a}$ is subdirect for every $a$. If $R_{a}$ is linked, then $S_{a}$ (defined analogously to $R_{a}$) is full. If $R_{a}$ is not linked, then it is the graph of a bijection $A\to A$ (since is simple), so $S_{a}$ is the equality relation. In summary, $S$ is equal to the relation $B(x)\vee(y=z)$. By Proposition 7.3, $B$ is strongly projective.
If $B$ is empty, then every $R_{a}$ is the graph of a bijection. But then, by symmetry of $R$, fixing any coordinate to any $a$ and projecting out this coordinate gives the graph of a bijection. Therefore $R$ is strongly functional, so is abelian by Proposition 4.8, finishing the proof.
We are ready to prove the main tool for Theorem 5.21. It will be useful to generalize the concept of a $B$-essential relation introduced in Subsection 7.5 to infinite powers: A subuniverse $R$ of is $B$-essential if $R$ does not intersect $B^{X}$ but every projection of $R$ onto all but one of the coordinates intersects the corresponding power of $B$. By Proposition 7.2, $\{a\}$ absorbs by a term operation of arity $n$ if and only if it does not have any $\{a\}$-essential subpower of arity $n$.
@@
Let be a Taylor algebra and $a\in A$. Suppose that $\{a\}$ does not absorb but does absorb every proper subalgebra of that contains $\{a\}$. Then
  • has a nontrivial abelian quotient or
  • there exists a nonempty projective subuniverse $B$ of such that $a\not\in B$.
Proof 9.36.
@@
First, observe that it is enough to prove the claim for simple . Indeed, if $\alpha$ is a maximal congruence of , then either $a/\alpha$ absorbs $A$ (in which case $\{a\}$ absorbs by transitivity of absorption, a contradiction to the assumptions), or it does not (in which case we apply the simple case – observe that projective subuniverses lift from quotients). Assume therefore that is simple.
@@
We further assume that every relation $S\leq_{sd}$ whose left center contains $a$ is full. Indeed, if it is not, then by Corollary 8.8 either has a proper projective and 2-absorbing subuniverse $B$ or the left center of $R$ is an absorbing subuniverse of . By transitivity of absorption and the assumptions, no absorbing subuniverse of can contain $a$. Therefore $B$ would be a projective subuniverse not containing $a$ and the proof would be concluded.
@@
Next we observe that there exists a symmetric $\{a\}$-essential relation $R\leq$ (here it will be convenient to use an infinitary relation). Indeed, for any arity $n$ there exists by Proposition 7.2 some $\{a\}$-essential relation, i.e., containing tuples $(b_{1},a,a,\dots,a)$, $(a,b_{2},a\dots,a),\dots$ and not containing $(a,a,\dots,a)$. Taking an element $b=b(n)$ that appears the most often among the $b_{i}$ and fixing (and projecting out) the other coordinates to $a$ we get an $\{a\}$-essential relation of arity at least $n/|A|$ containing $(b,a,a\dots)$, $(a,b,a,\dots)$, …. Finally, by taking $R=\operatorname{Sg}((b,a,a,\dots,),(a,b,a,\dots),\dots)$, where $b$ is an element that appears infinitely many times among the $b(n)$, we get a symmetric $\{a\}$-essential relation.
@@
For a symmetric relation $R\leq$, denote $B(R)=\{b\in A:(b,a,a,\dots)\in R\}$, and take a symmetric $\{a\}$-essential relation $R$ (i.e., $a\not\in B(R)\neq\emptyset$) such that $B(R)$ is maximal. Set $B=B(R)$. We will show that $R$ either gives us a ternary subpower forcing strong projectivity or abelianess, or $R$ is a witness for $B$ being a projective subuniverse, i.e., after fixing any co-finite collection coordinates to $a$ and projecting them out, $R$ becomes $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{k})$ (see Proposition 2.6); observe here that the obtained relation is indeed a subpower of despite the infinite arity of $R$. Since $a\not\in B$ we will be done.
@@
We claim that $\operatorname{Sg}(a,b)=A$ for every $b\in B$. Indeed, otherwise the relation $R$ defined by $\operatorname{Sg}\{(b,a,a,\dots)$, $(a,b,a,\dots),\dots\}$ is an $\{a\}$-essential symmetric subuniverse of $(\operatorname{Sg}(a,b))^{\mathbb{N}}$, so $\{a\}$ does not absorb $\operatorname{Sg}(a,b)$ by Proposition 7.2 (consider, as above, the relations obtained by fixing co-finite collections of coordinates to $a$), contradicting the assumptions of the theorem.
@@
If $B^{\prime}(=B+S+S)=B$, then $S$ is not linked, so it is a graph of a bijection by simplicity of . We fix all but three arbitrarily selected coordinates of $R$ to $a$, project onto the three coordinates, and call $T$ the obtained subuniverse of . The relation $T$ has full projection to the first two coordinates (as we argued above) and the binary relation obtained by fixing a coordinate to $a$ is the graph of a bijection. Lemma 9.33 now implies that either is abelian or has a nonempty strongly projective subuniverse that does not contain $a$. Since strong projectivity implies projectivity, we are done.